# Segment Anything Meets Point Tracking

### v1

@@ Abstract
The Segment Anything Model (SAM) has established itself as a powerful zero-shot image segmentation model, employing interactive prompts such as points to generate masks. This paper presents SAM-PT, a method extending SAM’s capability to tracking and segmenting anything in dynamic videos. SAM-PT leverages robust and sparse point selection and propagation techniques for mask generation, demonstrating that a SAM-based segmentation tracker can yield strong zero-shot performance across popular video object segmentation benchmarks, including DAVIS, YouTube-VOS, and MOSE. Compared to traditional object-centric mask propagation strategies, we uniquely use point propagation to exploit local structure information that is agnostic to object semantics. We highlight the merits of point-based tracking through direct evaluation on the zero-shot open-world Unidentified Video Objects (UVO) benchmark. To further enhance our approach, we utilize K-Medoids clustering for point initialization and track both positive and negative points to clearly distinguish the target object. We also employ multiple mask decoding passes for mask refinement and devise a point re-initialization strategy to improve tracking accuracy. Our code integrates different point trackers and video segmentation benchmarks and will be released at https://github.com/SysCV/sam-pt.
@@
Figure 1: Segment Anything Meets Point Tracking (SAM-PT). SAM-PT is the first method to utilize sparse point propagation for Video Object Segmentation (VOS). The essence of SAM-PT is to extend SAM [16] with long-term point trackers to effectively operate on videos in a zero-shot manner. SAM-PT takes a video as input together with annotations of the target object in the first frame. These annotations are called “query points” and denote either the target object (positive points) or designate non-target segments (negative points). The points are tracked throughout the video using point trackers that propagate the query points to all video frames, producing predicted trajectories and occlusion scores. SAM is subsequently prompted with the non-occluded points in the trajectories as to output a segmentation mask for each video frame independently.
@@ 1 Introduction
Video segmentation benefits a myriad of applications, including autonomous driving, robotics, and video editing. Despite significant progress made in the past few years with deep neural networks [41, 4, 37, 3], the current methodologies falter when faced with unseen data, particularly in zero-shot settings. These models struggle to maintain consistent performance across diverse scenarios without specific video segmentation data for fine-tuning.
@@
The prevailing methods [4, 3] in semi-supervised Video Object Segmentation (VOS) and Video Instance Segmentation (VIS) exhibit performance gaps when dealing with unseen data, particularly in a zero-shot setting, i.e., when these models are transferred to video domains they have not been trained on and encompass object categories that fall outside of the training distribution.
A potential route towards overcoming these challenges lies in adapting successful models in the image segmentation domain for video segmentation tasks. One such promising model is the Segment Anything Model (SAM) [16]. SAM is a powerful foundation model for image segmentation, trained on the large-scale SA-1B dataset, which contains an astounding 11 million images and over 1 billion masks. This extensive training set enables SAM’s impressive zero-shot generalization capabilities. The model is highly adaptable, able to produce high-quality masks from single foreground points, and has demonstrated robust performance across a range of downstream tasks under zero-shot transfer protocols. While SAM demonstrates powerful zero-shot capabilities for image segmentation, it is not innately suited for video segmentation tasks.
Recent efforts have been made to adapt SAM for video segmentation. For instance, TAM [40] integrates SAM with the state-of-the-art memory-based mask tracker XMem [4]. Likewise, SAM-Track [6] combines SAM with DeAOT [41]. While these methods mostly recover the performance on in-distribution data, they fall short in preserving the original performance of SAM in more challenging, zero-shot settings. Other methods that do not leverage SAM, such as SegGPT [36], can successfully solve a number of segmentation problems using visual prompting, but still require mask annotation for the first video frame. This problem represents a significant barrier in zero-shot video segmentation, particularly as we seek to develop methods that can easily generalize to unseen scenarios and consistently deliver high-quality segmentation across diverse video domains.
@@
We introduce SAM-PT (Segment Anything Meets Point Tracking), depicted in Segment Anything Meets Point Tracking. This is the first method to utilize sparse point tracking combined with SAM for video segmentation, offering a new perspective on solving the problem. Instead of employing object-centric dense feature matching or mask propagation, we propose a point-driven approach that capitalizes on tracking points using rich local structure information embedded in videos. As a result, it only requires sparse points annotation to denote target object in the first frame and provides better generalization to unseen objects, a strength demonstrated on the open-world UVO [34] benchmark. This approach also helps preserve the inherent flexibility of SAM while extending its capabilities effectively to video segmentation.
@@
SAM-PT prompts SAM with sparse point trajectories predicted using state-of-the-art point trackers, such as PIPS [11], harnessing their versatility for video segmentation. We identify that initializing points to track using K-Medoids cluster centers from a mask label was the strategy most compatible with prompting SAM. Tracking both positive and negative points enables the clear delineation of target objects from their background. To further refine the output masks, we propose multiple mask decoding passes that integrate both types of points. In addition, we devised a point re-initialization strategy that increases tracking accuracy over time. This approach involves discarding points that have become unreliable or occluded, and adding points from object parts or segments that become visible in later frames, such as when the object rotates.
Notably, our experimental results highlight that SAM-PT competes with existing zero-shot methods [36] or outperforms them [45, 2, 12, 35] on several video segmentation benchmarks. This comes without the need for any video segmentation data during training, underscoring the robustness and adaptability of our approach. SAM-PT holds the potential to enhance progress in video segmentation tasks, particularly in zero-shot scenarios.
@@ 2 Related Work
Point Tracking for Video Segmentation.
Classical feature extraction and tracking methods such as Lucas-Kanade [23], Tomasi-Kanade [31], Shi-Tomasi [29], SIFT [22], and SURF [1], as well as newer methods such as LIFT [43], SuperPoint [7], and SuperGlue [28], have all demonstrated proficiency in identifying or tracking sparse features and establishing long-range correspondences. However, their effectiveness is confined to a specific set of distinct interest points and they often struggle when applied to non-rigid, dynamic scenes. Flow-based methods, such as RAFT [30], excel in tracking dense points between successive frames. However, they stumble with deriving accurate long-range point trajectories. When chaining flow predictions over time, errors tend to accumulate and lead to drift, while occlusions result in tracking failures.
Significant strides have recently been made in long-term point tracking across video frames, as evinced by methods such as TapNet [9] and PIPS [11], as well as the concurrent and state-of-the-art OmniMotion [32] and TAPIR [10] techniques. These approaches optimize long-range point trajectories across an entire video, navigating mostly well through periods of occlusion.
Our work stands apart as the first to integrate these successful long-term point tracking methods, utilizing them to guide a promptable foundation model for image segmentation toward performing video segmentation tasks.
Segment and Track Anything models.
SAM [16] is an innovative image segmentation model for promptable image segmentation, trained on over 1 billion segmentation masks. It showcases remarkable zero-shot generalization abilities and can produce high-quality masks from a single foreground point. To further improve the quality of the masks, especially when segmenting objects with intricate structures, HQ-SAM [15] extends SAM with a learnable high-quality output token which proves efficient in diverse segmentation domains. However, SAM and HQ-SAM cannot be directly used to solve video segmentation tasks.
A few concurrent works extend SAM, for example, TAM [40] and SAM-Track [6] combine SAM with state-of-the-art mask trackers (such as XMem [4] and DeAOT [42]) to perform interactive video object segmentation. These methods employ SAM for mask initialization or correction and XMem/DeAOT for mask tracking and prediction. Using the pre-trained mask trackers recovers the in-distribution performance, but hinders the performance in zero-shot settings. PerSAM [45] also demonstrates the ability to track multiple reference objects in a video. Instead of building an interactive tracking pipeline or SAM fine-tuning, we focus on learning robust associations for diverse objects in zero-shot scenarios.
Zero-shot VOS / VIS.
Among the non-SAM-based methods, Painter [35] and its SegGPT [36] extension are another sort of generalist models for solving a variety of image and segmentation tasks. These methods likewise use visual prompting techniques but are inherently different frameworks from SAM. Despite its wide applicability, Painter shows lacking performance in video segmentation tasks. Conversely, SegGPT successfully uses in-context prompting to achieve one-shot video object segmentation performance comparable to ours, also without training on any video data. The training domains, however, notably differ between SegGPT and our method.
STC [12] and DINO [2] also do not use any video segmentation data during training. In the semi-supervised video object segmentation, they take a reference mask as input and perform frame-by-frame feature matching, which propagates the reference mask across the entirety of the video. Our SAM-PT, on the other hand, diverges substantially from these methodologies by adopting point tracking, eschewing the process of frame-by-frame feature matching. Additionally, our method requires only sparse points to represent the target object, rather than a full reference mask, and yields superior performance on conventional semi-supervised video object segmentation benchmarks.
@@ 3 Method
We propose SAM-PT to adapt SAM, a foundation model for image segmentation, for addressing video segmentation tasks in a zero-shot setting. SAM-PT combines the strengths of existing prominent point trackers, such as PIPS [11] and TapNet [9], with the powerful image segmentation of SAM to enable tracking of anything in videos. First, Sec. 3.1 briefly describes SAM. Sec. 3.2 then introduces our SAM-PT method with its four constituent steps. Finally, Sec. 3.3 analyzes and highlights the method’s novelty as the first point-driven video segmentation method compared to existing works.
@@ 3.1 Preliminaries: SAM
The Segment Anything Model (SAM) [16] is a novel vision foundation model designed for promptable image segmentation. SAM is trained on the large-scale SA-1B dataset, which contains 11 million images and over 1 billion masks. SA-1B has 400 times more masks than any existing segmentation dataset. This extensive training set facilitates SAM’s impressive zero-shot generalization capabilities to new data. SAM has showcased its ability to produce high-quality masks from a single foreground point, and has demonstrated robust generalization capacity on a variety of downstream tasks under a zero-shot transfer protocol using prompt engineering. These tasks include, but are not limited to, edge detection, object proposal generation, and instance segmentation.
SAM comprises of three main components: an image encoder, a flexible prompt encoder, and a fast mask decoder. The image encoder is a Vision Transformer (ViT) backbone and processes high-resolution $1024\times 1024$ images to generate an image embedding of $64\times 64$ spatial size. The prompt encoder takes sparse prompts as input, including points, boxes, and text, or dense prompts such as masks, and translates these prompts into $c$-dimensional tokens. The lightweight mask decoder then integrates the image and prompt embeddings to predict segmentation masks in real-time, allowing SAM to adapt to diverse prompts with minimal computational overhead.
@@ 3.2 Ours: SAM-PT
Figure 2: Segment Anything Meets Point Tracking (SAM-PT) overview. The essence of SAM-PT is to extend image segmentation foundation model to effectively operate on videos. SAM-PT has four steps: $\mathbf{1}$) Query Points Selection. It starts with first-frame query points which denote the target object (positive points) or designate non-target segments (negative points). These points are provided by the user or derived from a ground truth mask. $\mathbf{2}$) Point Tracking. Initiated with the query points, our approach leverages point trackers to propagate the points across video frames, predicting point trajectories and occlusion scores. $\mathbf{3}$) Segmentation. The trajectories are then used to prompt the Segment Anything Model (SAM) and output per-frame mask predictions. $\mathbf{4}$) Point Tracking Reinitialization. Optionally, the predicted masks are used to reinitialize the query points and restart the process when reaching a prediction horizon $h$. Re-initialization helps by getting rid of unreliable and occluded points, and adds points from object parts or segments that become visible in later frames, such as when the object rotates.
1) Query Points Selection.
The process begins with defining query points in the first video frame, which either denote the target object (positive points) or designate the background and non-target objects (negative points). Users can manually and interactively provide query points, or they may be derived from a ground truth mask. For example, in the case of semi-supervised video object segmentation, the ground truth mask is provided for the first frame where the object appears. We derive the query points from ground truth masks using different point sampling techniques by considering their geometrical locations or feature dissimilarities, as depicted in Fig. 3. These sampling techniques are:
@@
While each method contributes distinct characteristics that influence the model’s performance, our ablation study reveals that K-Medoids sampling yields the best results with a good full coverage of various segments of the complete object. Shi-Tomasi sampling follows closely, indicating their respective strengths in this context. The selection and arrangement of these points considerably affect the overall video segmentation performance, thus determining the optimal method is crucial.
@@
2) Point Tracking.
Initiated with the query points, we employ robust point trackers to propagate the points across all frames in the video, resulting in point trajectories and occlusion scores. We adopt the state-of-the-art point tracker PIPS [11] to propagate the points as PIPS shows moderate robustness toward long-term tracking challenges such as object occlusion and re-appearance. This is also shown more effective than methods such as chained optical flow propagation or first-frame correspondences in our experiment section.
3) Segmentation.
In the predicted trajectories, the non-occluded points serve as indicators of where the target object is throughout the video. This allows us to use the non-occluded points to prompt SAM, as illustrated in Fig. 4, and leverage its inherent generalization ability to output per-frame segmentation mask predictions. Unlike conventional tracking methods that require training or fine-tuning on video segmentation data, our approach excels in zero-shot video segmentation tasks.
@@
4) Point Tracking Reinitialization.
We optionally execute a reinitialization of the query points using the predicted masks once a prediction horizon of $h=8$ frames is reached, and denote the variant as SAM-PT-reinit. Upon reaching this horizon, we have $h$ predicted masks and will take the last predicted mask to sample new points. At this stage, all previous points are discarded and substituted with the newly sampled points. Following this, steps 1) through 4) are repeated with the new points, starting from the horizon timestep where reinitialization occurs. The steps are iteratively executed until the entire video is processed. The reinitialization process serves to enhance tracking accuracy over time by discarding points that have become unreliable or occluded, while incorporating points from object segments that become visible later in the video. Other reinitialization variants are discussed in Appendix A and included in the ablation study in Sec. 4.3.
@@
With sparse point tracking combined with prompting SAM, SAM-PT distinguishes itself from traditional video segmentation methods that depend on dense object mask propagation, as noted in Tab. 1. To propagate the first-frame GT label to the remaining video frames, traditional techniques commonly use feature matching with masks cached to a mask memory [40, 6, 4, 41], frame-by-frame feature matching [12, 2], feature matching with the first-frame mask [45], optical flow [39], and, recently, in-context visual prompting [35, 36]. In contrast, SAM-PT introduces a unique approach to video object segmentation, employing the robust combination of point tracking with SAM, which is inherently designed to operate on sparse point prompts.
@@
In summary, SAM-PT is the first method that introduces sparse point propagation combined with prompting a image segmentation foundation model to perform zero-shot video object segmentation. It provides a fresh perspective and adds a new dimension to the study of video object segmentation.
Table 1: Comparison of semi-supervised Video Object Segmentation (VOS) methods with respect to mask annotation requirements and propagation techniques. Our method, SAM-PT, is the first method for VOS that uses sparse point propagation. With such compact mask representation, we achieve the highest $\mathcal{J\&F}$ scores on the DAVIS $2016$ and $2017$ validation subsets among methods that do not utilize any video segmentation data during training. The methods are compared based on their use of video mask data during training, whether they are evaluated zero-shot on DAVIS, what first-frame labels they require, and what label propagation technique they employ.
Method Video Mask Zero- Shot Frame Init. Propagation DAVIS $\mathbf{2016}$ DAVIS $\mathbf{2017}$
@@
XMem [4] Mask Feature Matching $92.0$ 87.7
DeAOT [41] Mask Feature Matching $\mathbf{92.9}$ $86.2$
Painter [35] Mask Mask Prompting - $34.6$
@@
PerSAM-F [45] Mask Feature Matching - $71.9$
SegGPT [36] Mask Mask Prompting $82.3$ $75.6$
SAM-PT (ours) Points Points Prompting $\mathbf{83.1}$ $\mathbf{76.6}$

### v2

@@ Abstract
The Segment Anything Model (SAM) has established itself as a powerful zero-shot image segmentation model, enabled by efficient point-centric annotation and prompt-based models. While click and brush interactions are both well explored in interactive image segmentation, the existing methods on videos focus on mask annotation and propagation. This paper presents SAM-PT, a novel method for point-centric interactive video segmentation, empowered by SAM and long-term point tracking. SAM-PT leverages robust and sparse point selection and propagation techniques for mask generation. Compared to traditional object-centric mask propagation strategies, we uniquely use point propagation to exploit local structure information agnostic to object semantics. We highlight the merits of point-based tracking through direct evaluation on the zero-shot open-world Unidentified Video Objects (UVO) benchmark. Our experiments on popular video object segmentation and multi-object segmentation tracking benchmarks, including DAVIS, YouTube-VOS, and BDD100K, suggest that a point-based segmentation tracker yields better zero-shot performance and efficient interactions. We release our code that integrates different point trackers and video segmentation benchmarks at https://github.com/SysCV/sam-pt.
@@
Figure 1: Segment Anything Meets Point Tracking (SAM-PT). SAM-PT is a point-centric method that utilizes sparse point propagation for interactive video segmentation, enabling easier interaction and faster annotation. We extend SAM [21] with long-term point trackers to effectively operate on videos in a zero-shot manner. SAM-PT takes user clicks as “query points” which either denote the target object (positive points) or designate non-target segments (negative points). The points are tracked throughout the video using point trackers that propagate the query points to all video frames, producing trajectory predictions and occlusion scores. SAM is subsequently prompted with the non-occluded points in the trajectories as to output a segmentation mask for each video frame independently. The propagated points can be further edited for accurate segmentation and tracking.
@@ 1 Introduction
Object segmentation and tracking in videos are central pillars for a myriad of applications, including autonomous driving, robotics, and video editing. Despite significant progress made in the past few years with deep neural networks [48, 7, 43, 5], we still need to rely on expensive labels for model supervision to achieve high accuracies. Therefore, many efforts have been made on interactive video segmentation to accelerate the data labeling process and benefit artistic video editing. Those methods are usually evaluated on a simplified semi-supervised segmentation setup as a proxy for full interactive video segmentation.
@@
The prevailing methods [7, 5] in semi-supervised Video Object Segmentation (VOS) and Video Instance Segmentation (VIS) exhibit performance gaps when dealing with unseen data, particularly in a zero-shot setting, i.e., when these models are transferred to video domains they have not been trained or that encompass object categories falling outside of the training distribution.
Further, generalizable models usually require large amounts of training data. The existing interactive video segmentation methods assume the mask of an object is given on the first frame of the testing video. While getting accurate mask is laborious, recent works [21, 6] on training foundation image segmentation models show that point-based annotation in combination with mask editing tools is a scalable approach to label exceedingly large amounts of data. Despite its success on images in terms of labeling efficiency and accuracy, point-centric interactive segmentation has received scant attention in the video domains.
In this paper, we aim to achieve both domain generalizability and labeling efficiency for interactive video segmentation. Our insights are two-fold. First, foundation models in image segmentation are available, such as Segment Anything Model (SAM) [21]. SAM, trained on 11 million images and 1 billion object masks, has impressive zero-shot generalization capabilities. The model also supports point prompts as additional inputs for interactive image segmentation and produces high-quality masks. Second, we witnessed significant recent progress in point tracking [14, 16, 53, 15, 39, 19]. Those tracking methods, once trained, can propagate points across video frames on diverse domains.
@@
Therefore, we introduce SAM-PT (Segment Anything Meets Point Tracking), depicted in Segment Anything Meets Point Tracking. This is the first method to utilize sparse point tracking combined with SAM for video segmentation, offering a new perspective on solving the problem. Instead of employing object-centric dense feature matching or mask propagation, we propose a point-centric approach that capitalizes on tracking points using rich local structure information embedded in videos. It only requires sparse points annotation to denote the target object in the first frame and provides better generalization to unseen objects. This approach also helps preserve the inherent flexibility of SAM while extending its capabilities effectively to video segmentation. Similar to the data annotation process in SAM, our point-centric approach can be potentially integrated with the existing mask-based approaches in real-world applications.
@@
SAM-PT prompts SAM with sparse point trajectories predicted using state-of-the-art point trackers, such as CoTracker [19], harnessing their versatility for video segmentation. We identified that initializing points to track using K-Medoids cluster centers from a mask label was the strategy most compatible with prompting SAM. Tracking both positive and negative points enables the clear delineation of target objects from their background. To further refine the output masks, we propose multiple mask decoding passes that integrate both types of points. In addition, we devised a point reinitialization strategy that increases tracking accuracy over time. This approach involves discarding points that have become unreliable or occluded, and adding points from object parts or segments that become visible in later frames, such as when the object rotates.
We evaluate SAM-PT on multiple setups including semi-supervised, open-world, and fully interactive video segmentation. Our method achieves stronger performance than existing zero-shot methods by up to $5.0\%$ on DAVIS, $2.0\%$ on YouTube-VOS, and $7.3\%$ on BDD100K, while also surpassing a fully-supervised VIS method [46] on UVO by $6.7$ points. We also set up a new benchmark for interactive point-based video segmentation to simulate the process of manually labeling the whole video. In this setup, SAM-PT significantly reduces annotation effort, approaching the performance of fully supervised approaches and underscoring its practicality. This comes without the need for any video segmentation data during training, underscoring the robustness and adaptability of our approach, and indicating its potential to enhance progress in video segmentation tasks, particularly in zero-shot scenarios.
@@ 2 Related Work
Point Tracking for Video Segmentation. Classical feature extraction and tracking methods such as Lucas-Kanade [26], Tomasi-Kanade [36], Shi-Tomasi [34], SIFT [25], and SURF [1], as well as newer methods such as LIFT [49], SuperPoint [12], and SuperGlue [33], have all demonstrated proficiency in identifying or tracking sparse features and establishing long-range correspondences. Nonetheless, these techniques often falter in dynamic, non-rigid environments. While flow-based approaches such as RAFT [35] offer improvements, they too struggle with maintaining long-term point accuracy due to error accumulation and occlusions. Addressing these shortcomings, recent innovations such as PIPS [16], PIPS++ [53], OmniMotion [39], TAPIR [15], and the state-of-the-art CoTracker [19], optimize for robust long-term trajectories and effectively manage occlusions. Our work is unique in applying these methods to guide image segmentation models for video segmentation tasks.
Segment and Track Anything Models. SAM [21] is a foundation model for image segmentation that showcases impressive zero-shot capabilities. Its extension, HQ-SAM [20], improves mask quality for complex objects but is not designed for video tasks. TAM [46] and SAM-Track [11] attempt to extend SAM to video segmentation by integrating the state-of-the-art fully-supervised XMem [7] and DeAOT [47] mask trackers, respectively, yet they lack in zero-shot scenarios.
Zero-Shot VOS / VIS. Generalist models such as Painter [41] apply visual prompting to various tasks but demonstrate limited performance in video segmentation. On the other hand, SegGPT [42] also uses visual prompting and competes closely with our method on some datasets. Other approaches, such as STC [18] and DINO [4], perform VOS through feature matching. Our approach distinguishes itself by taking the point-centric approach to enhance performance on VOS benchmarks in a zero-shot setting.
Interactive VOS. Interactive VOS has shifted from labor-intensive manual annotations to more user-friendly interaction methods, such as scribbles, clicks, and drawings, enabling rapid and intuitive video editing [37, 28, 27, 3, 17]. Among these, MiVOS [8] stands out for its modular design that decouples mask generation from propagation, effectively incorporating user interactions from diverse interaction modalities. Unlike MiVOS and other fully-supervised methods, SAM-PT is the first to use point propagation instead of mask propagation and thus operates effectively in zero-shot settings. Our interactive point-based video segmentation study emphasizes the simplicity and efficacy of point interactions and differs from common scribble-based benchmarking [2] or in-distribution user studies [8].
@@ 3 Method
We propose SAM-PT for addressing video segmentation tasks in a zero-shot setting. SAM-PT combines the strengths of the Segment Anything Model (SAM), a foundation model for image segmentation, and prominent point trackers, such as PIPS [16] and CoTracker [19], to enable interactive tracking of anything in videos. Sec. 3.1 briefly describes the background knowledge about SAM. Sec. 3.2 then introduces our SAM-PT method with its four constituent steps. Finally, Sec. 3.3 analyzes and highlights the method’s novelty as the first point-centric interactive video segmentation method compared to existing works.
@@ 3.1 Preliminaries: SAM
Whereas in computer vision “zero-shot (learning)” usually refers to the study of generalization to unseen object categories in image classification [22], we follow prior work [21, 31] and rather employ the term in a broader sense and explore generalization to unseen datasets.
The Segment Anything Model (SAM) [21] is a novel vision foundation model designed for promptable image segmentation. SAM is trained on the large-scale SA-1B dataset, which contains 11 million images and over 1 billion masks. SA-1B has 400 times more masks than any prior segmentation dataset. This extensive training set facilitates SAM’s impressive zero-shot generalization capabilities. SAM has showcased its ability to produce high-quality masks from a single foreground point and has demonstrated robust generalization capacity on a variety of downstream tasks under a zero-shot transfer protocol using prompt engineering. These tasks include, but are not limited to, edge detection, object proposal generation, and instance segmentation.
SAM comprises three main components: an image encoder, a flexible prompt encoder, and a fast mask decoder. The image encoder is a Vision Transformer (ViT) backbone and processes high-resolution $1024\times 1024$ images to generate an image embedding of $64\times 64$ spatial size. The prompt encoder takes sparse prompts as input, including points, boxes, and text, or dense prompts such as masks, and translates these prompts into $c$-dimensional tokens. The lightweight mask decoder then integrates the image and prompt embeddings to predict segmentation masks in real-time, allowing SAM to adapt to diverse prompts with minimal computational overhead.
@@ 3.2 Ours: SAM-PT
Figure 2: Segment Anything Meets Point Tracking (SAM-PT) overview. The essence of SAM-PT is to extend image segmentation foundation models to effectively operate on videos. SAM-PT has four steps: $\mathbf{1}$) Query Points Selection. It starts with first-frame query points which denote the target object (positive points) or designate non-target segments (negative points). These points are provided by the user or derived from a ground truth mask. $\mathbf{2}$) Point Tracking. Initiated with the query points, our approach leverages point trackers to propagate the points across video frames, predicting point trajectories and occlusion scores. $\mathbf{3}$) Segmentation. The trajectories are then used to prompt the Segment Anything Model (SAM) and output per-frame mask predictions. $\mathbf{4}$) Point Tracking Reinitialization. Optionally, the predicted masks are used to reinitialize the query points and restart the process when reaching a prediction horizon $h$. Reinitialization helps by getting rid of unreliable points and adding points to object segments that become visible in later frames.
1) Query Points Selection. The process begins with defining query points in the first video frame, which either denote the target object (positive points) or designate the background and non-target objects (negative points). Users can manually and interactively provide query points, or they may be derived from a ground truth mask. For example, in the case of semi-supervised video object segmentation, the ground truth mask is provided for the first frame where the object appears. We derive the query points from ground truth masks using different point sampling techniques by considering their geometrical locations or feature dissimilarities, as depicted in Fig. 3. These sampling techniques are:
@@
While each method contributes distinct characteristics that influence the model’s performance, our ablation study reveals that K-Medoids sampling yields the best results with good coverage of various segments of the complete object. Shi-Tomasi sampling follows closely, indicating their respective strengths in this context. The selection and arrangement of these points considerably affect the overall video segmentation performance, thus determining the optimal method is crucial.
@@
2) Point Tracking. Initiated with the query points, we employ robust point trackers to propagate the points across all frames in the video, resulting in point trajectories and occlusion scores. We adopt point trackers such as PIPS [16] and the state-of-the-art CoTracker [19] to propagate the points as they show moderate robustness toward long-term tracking challenges such as object occlusion and re-appearance. Long-term point trackers are also shown more effective than methods such as chained optical flow propagation or first-frame correspondences in our experiments.
3) Segmentation. In the predicted trajectories, the non-occluded points serve as indicators of where the target object is throughout the video. This allows us to use the non-occluded points to prompt SAM, as illustrated in Fig. 4, and leverage its inherent generalization ability to output per-frame segmentation mask predictions. Unlike conventional tracking methods that require training or fine-tuning on video segmentation data, our approach excels in zero-shot video segmentation tasks.
@@
4) Point Tracking Reinitialization. We optionally execute a reinitialization of the query points using the predicted masks once a prediction horizon of $h=8$ frames is reached. Upon reaching this horizon, we have $h$ predicted masks and will take the last one to sample new points. At this stage, all previous points are discarded and substituted with the newly sampled points. Following this, steps 1) through 4) are repeated with the new points, starting from the horizon timestep where reinitialization occurs. The steps are iteratively executed until the entire video is processed. The reinitialization process serves to enhance tracking accuracy over time by discarding unreliable or occluded points while incorporating points from object segments that become visible later in the video. Other reinitialization variants are discussed in our Supplementary Material.
@@
With sparse point tracking combined with prompting SAM, SAM-PT distinguishes itself from traditional video segmentation methods that depend on dense object mask propagation, as noted in Tab. 1. To propagate the first-frame GT label to the remaining video frames, traditional techniques commonly use feature matching with masks cached to a mask memory [46, 11, 7, 48], frame-by-frame feature matching [18, 4], optical flow [45], and, recently, in-context visual prompting [41, 42]. In contrast, SAM-PT introduces a unique approach to video object segmentation, employing the robust combination of point tracking with SAM, which is inherently designed to operate on sparse point prompts.
@@
Table 1: Comparative analysis of semi-supervised Video Object Segmentation methods. Our approach, SAM-PT, introduces sparse point propagation, a compact mask representation that uses local structure information agnostic to object semantics. It outperforms other non-video-data-dependent methods, achieving top $\mathcal{J\&F}$ scores on DAVIS $2016$ and $2017$, and the highest $\mathcal{G}$ score on YouTube-VOS $2018$. The comparison considers the reliance on video mask data during training, zero-shot learning setting, initial frame label requirements, and label propagation techniques used.
Method Video Mask Zero- Shot Frame Init. Propagation DAVIS $\mathbf{2016}$ DAVIS $\mathbf{2017}$ YTVOS $\mathbf{2018}$
@@
DEVA [10] Mask Feature Matching - $87.6$
XMem [7] Mask Feature Matching $92.0$ 87.7 $86.1$
DeAOT [48] Mask Feature Matching $\mathbf{92.9}$ $86.2$ $\mathbf{86.2}$
Painter [41] Mask Mask Prompting - $34.6$ $24.1$
@@
SegGPT [42] Mask Mask Prompting $82.3$ $75.6$ $74.7$
SAM-PT (ours) Points Points Prompting $\mathbf{84.3}$ $\mathbf{79.4}$ $\mathbf{76.2}$